OpenStudy (anonymous):

Determine the domain of the function. f(x)= sqrt of 9-x

OpenStudy (cybershadow):

domain , eh well as the function is f(x) = Sqrt(9-x) therefore 9-x >= 0 (otherwise it would be an imaginary number) and solve the inequality that will give -x >=9 (multiplying by negative sign , and changing the inequality) x <= -9 so domain is x belongs to (-infinity,-9]

OpenStudy (anonymous):

so would it be all real numbers?

OpenStudy (anonymous):

ohh x is less than or equal to -9?

OpenStudy (cybershadow):

YEs!

OpenStudy (anonymous):

x ≤ 9 All real numbers except 9 All real numbers x > 9

OpenStudy (anonymous):

but those are my only options

OpenStudy (anonymous):

there is no negative 9

OpenStudy (cybershadow):

oh i think i made a type during solving the inequality 9-x <=0 -x <=-9 (multiplying by negative sign) x <= 9

OpenStudy (anonymous):

thank you ! (: